# Surface solar radiation: AI satellite retrieval can outperform Heliosat and generalizes well to other climate zones

### v1

@@ 7 Acknowledgements
This work was supported by the Swiss Innovation Agency Innosuisse through grant 104.531 IP-EE. We acknowledge helpful discussions with Jörg Trentmann and Uwe Pfeifroth of DWD. We made use of the pyranometer measurements from IEA-PVPS that are accessible via [50]. The DWD and KNMI pyranometer measurements are available from DWD [51] and KNMI [52]. The MeteoSwiss pyranometer data were provided by the Swiss Federal Office of Meteorology MeteoSwiss via IDAWEB [53]. The digital elevation model was provided by Copernicus [42].
@@ Appendix B Cloudy and Clear-sky comparisons
(C) Finetuned ¿ MeteoSwiss |
(D) Finetuned ¿¿ IEA-PVPS |
@@
MBE |
Station |
SARAH-3 |
(A) Emulator |
(B) Finetuned DWD |
(C) Finetuned ¿ MeteoSwiss |
(D) Finetuned ¿¿ IEA-PVPS |
(E) Trained on DWD |
CAB | 2 | 5 | 5 | 22 | 16 | -10 | |
CAR | 11 | 3 | -11 | -12 | -7 | -32 | |
CEN | 12 | 9 | -13 | -1 | 1 | -25 | |
GHA | 21 | 31 | -30 | -27 | 2 | -101 | |
MIL | 15 | 8 | -8 | 2 | 1 | -12 | |
MIS | 5 | 15 | 7 | -25 | -4 | -60 | |
NOR | -6 | -5 | -10 | 1 | -5 | -9 | |
OUJ | 15 | 16 | -1 | 1 | -3 | -34 | |
PAL | 5 | 7 | -18 | -1 | -8 | -26 | |
PAY | 2 | 10 | 0 | 15 | 12 | -2 | |
TAB | 1 | -10 | -28 | -36 | -14 | -74 | |
TAT | 21 | 29 | -36 | -21 | -3 | -89 | |
TOR | -1 | 1 | 1 | 14 | 5 | 3 | |
All | VIS | -10 | -1 | 1 | 11 | -1 | 2 |
CAB | -28 | -38 | -22 | -9 | -8 | -37 | |
CAR | -8 | -22 | -27 | -33 | -23 | -48 | |
CEN | -19 | -32 | -42 | -37 | -29 | -53 | |
GHA | 4 | 19 | -43 | -41 | -7 | -116 | |
MIL | -5 | -20 | -28 | -23 | -19 | -28 | |
MIS | -20 | -7 | -8 | -50 | -22 | -80 | |
NOR | -37 | -45 | -43 | -33 | -37 | -38 | |
OUJ | -7 | -8 | -16 | -17 | -19 | -51 | |
PAL | -30 | -41 | -56 | -44 | -44 | -59 | |
PAY | -26 | -32 | -26 | -13 | -11 | -23 | |
TAB | -18 | -32 | -44 | -54 | -28 | -96 | |
TAT | 7 | 17 | -49 | -34 | -11 | -103 | |
TOR | -37 | -48 | -35 | -23 | -30 | -26 | |
Clear-sky | VIS | -35 | -34 | -19 | -7 | -21 | -14 |
CAB | 26 | 42 | 27 | 49 | 37 | 13 | |
CAR | 51 | 56 | 22 | 33 | 27 | 1 | |
CEN | 50 | 61 | 23 | 44 | 38 | 10 | |
GHA | 117 | 102 | 46 | 52 | 56 | -16 | |
MIL | 47 | 56 | 26 | 43 | 34 | 14 | |
MIS | 72 | 76 | 45 | 43 | 45 | -4 | |
NOR | 23 | 32 | 19 | 32 | 23 | 17 | |
OUJ | 74 | 84 | 40 | 50 | 43 | 16 | |
PAL | 36 | 49 | 15 | 36 | 24 | 3 | |
PAY | 34 | 57 | 30 | 47 | 38 | 21 | |
TAB | 65 | 63 | 26 | 25 | 31 | 3 | |
TAT | 90 | 89 | 25 | 36 | 34 | -23 | |
TOR | 26 | 37 | 27 | 42 | 32 | 25 | |
Cloudy | VIS | 17 | 36 | 23 | 32 | 21 | 20 |

### v2

@@ 7 Acknowledgements
We thank the anonymous reviewers for the constructive comments. This work was supported by the Swiss Innovation Agency Innosuisse through grant 104.531 IP-EE. We acknowledge helpful discussions with Jörg Trentmann and Uwe Pfeifroth of DWD. We made use of the pyranometer measurements from IEA-PVPS that are accessible via (Forstinger et al. 2023). The DWD and KNMI pyranometer measurements are available from DWD (2024) and KNMI (2024). The MeteoSwiss pyranometer data were provided by the Swiss Federal Office of Meteorology MeteoSwiss via IDAWEB (MeteoSwiss 2024). The digital elevation model was provided by Copernicus (AIRBUS 2024).
@@ Appendix B Cloudy and Clear-sky comparisons
Instantaneous Root Mean Square Error (RMSE) [] calculated on all IEA-PVPS stations in the dataset. Clear-sky and cloudy are defined as and CSI 0.8.
@@
(C) Finetuned MeteoSwiss |
(D) Finetuned IEA-PVPS |
@@
Instantaneous Mean Bias Error (MBE) [] calculated on all IEA-PVPS stations in the dataset. Clear-sky and cloudy are defined as CSI 0.8 and CSI 0.8. MBE Station SARAH-3 (A) Emulator (B) Finetuned DWD (C) Finetuned MeteoSwiss (D) Finetuned IEA-PVPS (E) Trained on DWD CAB 2 5 5 22 16 -10 CAR 11 3 -11 -12 -7 -32 CEN 12 9 -13 -1 1 -25 GHA 21 31 -30 -27 2 -101 MIL 15 8 -8 2 1 -12 MIS 5 15 7 -25 -4 -60 NOR -6 -5 -10 1 -5 -9 OUJ 15 16 -1 1 -3 -34 PAL 5 7 -18 -1 -8 -26 PAY 2 10 0 15 12 -2 TAB 1 -10 -28 -36 -14 -74 TAT 21 29 -36 -21 -3 -89 TOR -1 1 1 14 5 3 All VIS -10 -1 1 11 -1 2 CAB -28 -38 -22 -9 -8 -37 CAR -8 -22 -27 -33 -23 -48 CEN -19 -32 -42 -37 -29 -53 GHA 4 19 -43 -41 -7 -116 MIL -5 -20 -28 -23 -19 -28 MIS -20 -7 -8 -50 -22 -80 NOR -37 -45 -43 -33 -37 -38 OUJ -7 -8 -16 -17 -19 -51 PAL -30 -41 -56 -44 -44 -59 PAY -26 -32 -26 -13 -11 -23 TAB -18 -32 -44 -54 -28 -96 TAT 7 17 -49 -34 -11 -103 TOR -37 -48 -35 -23 -30 -26 Clear-sky VIS -35 -34 -19 -7 -21 -14 CAB 26 42 27 49 37 13 CAR 51 56 22 33 27 1 CEN 50 61 23 44 38 10 GHA 117 102 46 52 56 -16 MIL 47 56 26 43 34 14 MIS 72 76 45 43 45 -4 NOR 23 32 19 32 23 17 OUJ 74 84 40 50 43 16 PAL 36 49 15 36 24 3 PAY 34 57 30 47 38 21 TAB 65 63 26 25 31 3 TAT 90 89 25 36 34 -23 TOR 26 37 27 42 32 25 Cloudy VIS 17 36 23 32 21 20